OpenStudy (anonymous):

Solve the equation, using the domain of x = {-1, 0, 2}. Write your solutions as a set of ordered pairs. Do not type any spaces. Where needed, round to two decimal places. 2x + 3y = 7

OpenStudy (campbell_st):

substitute the domain values but 1st rewrite the equation \[y = \frac{1}{3}(7 - 2x)\] x = -1 y = 3 x = 0 y = 7/3 x = 2 y = 1 ordered pairs (-1,3) (0, 7/3), (2, 1)